Tiger Furnishings produces two models of cabinets for home theater components, the Basic and the Dominator. Data on operations and costs for March follow:
 


BasicDominatorTotal
Units produced
1,140

330

1,470
Machine-hours
3,400

3,600

7,000
Direct labor-hours
2,000

2,000

4,000
Direct materials costs$13,000
$4,750
$17,750
Direct labor costs
62,500

52,500

115,000
Manufacturing overhead costs






202,970
Total costs





$335,720

 
Tiger Furnishings’s CFO believes that a two-stage cost allocation system would give managers better cost information. She asks the company’s cost accountant to analyze the accounts and assign overhead costs to two pools: overhead related to direct labor cost and overhead related to machine-hours.

 

The analysis of overhead accounts by the cost accountant follows:

 

Manufacturing OverheadOverhead
Estimate
Cost Pool Assignment
Utilities$1,600Machine-hour related
Supplies
4,100Direct labor cost related
Training
8,200Direct labor cost related
Supervision
25,800Direct labor cost related
Machine depreciation
27,000Machine-hour related
Plant depreciation
27,400Machine-hour related
Miscellaneous
108,870Direct labor cost related


Required:

b. Compute the product costs per unit assuming that Tiger Furnishings uses direct labor costs and machine-hours to allocate overhead to the products. (Do not round intermediate calculations. Round your final answers to the nearest whole number.)

                                                basic                dominator                    total

product costing (blank row)

  direct material

  direct labor

Overhead (blank row)

  machine-related

  labor-related

   total overhead

total cost

units produced

unit cost
Answer #1

step 1) machine total overhead  56,000/7000= 8

           labor total overhead         146,970/115,000= 1.278


Basic product costing

DM= 13,000 from chart 1

DL= 62,500 from chart 1

basic overhead

machine related:  3400*8= 27,200

labor related: 62,500*1.278= 79,875

total overhead 27,200+79,875= 107,075

total cost: 107,025+ 13,000+ 62,500= 182,575

units produced: 1,140 from chart 1

unit cost: 182,575/1,140= 160.1535087719 or 160


dominator product

Dm= 4,750

dL= 52,500 

overhead

machine related 3600*8= 28,800

labor related 52,500*1.278=67,095

total overhead 28,800+67,095= 95,895

total cost: 95,895+4,750+52,500= 153,145

units produced 330

unit cost 153,145/330= 464.0757575758 or 464


total column: basic+ dominator all except units produced and unit cost.